Use the information in the following table to answer to 10 - Exports of goods and services Imports of goods and services Net not an economics major need help
Answer #1

(a) Trade balance = Exports- Imports

= (1000-800)= 200

(b) Net Unilateral transfer = (Unilateral transfers received - Unilateral transfers paid)

= (100-200) = -100

(c) Balance on the current account = (Trade balance + Net Unilateral transfer)

= (200+ (-100) ) = 100

(d) Balance on the financial account = (Net change in assets owned abroad - Net change in foreign owned assets owned at home )+ ( Investment income received from foreigners- Investment income paid to foreigners )

= (500-400) +(400-300)

= 100 +100 = 200

(e) Statistical discrepancy = Balance in financial account - Balance in current account

= (200-100 = 100